| TIP: Click on subject to list as thread! | ANSI |
| echo: | |
|---|---|
| to: | |
| from: | |
| date: | |
| subject: | 7\09 Pt 1 ISS On-Orbit Status 09-07-2003 |
This Echo is READ ONLY ! NO Un-Authorized Messages Please!
~~~~~~~~~~~~~~~~~~~~~~~~ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
10 July 2003
ISS On-Orbit Status 9 Jul 2003
Part 1 of 2
All ISS systems continue to function nominally, except as noted
previously or below. The crew awoke to kudos for yesterday's great
performance on the PAO/TV event with students and NASDA
Shuttle-Astronaut Dr. Mamoru Mohri assembled in Tokyo, Japan. ["You
made a lot of faces smile over there."]
CDR Yuri Malenchenko tackled a major 2-hr. IFM (in-flight
maintenance) on the Service Module (SM) condensate water recovery
system, removing and replacing the no. 2 condensate separation &
pumping unit (BRPK-2). [After deinstalling the old unit and tagging
up with TsUP (which deactivated power on the BITS2-12 on-board data
telemetry system), he disconnected the BITS and installed a
previously used BRPK plus a new pipe conduit delivered on Progress
11P, following instructions listed in an on-board IFM procedures data
file.]
FE/SO Dr. Ed Lu meanwhile spent time in the Lab working on U.S.
segment (USOS) computers, first rebuilding and restoring two Station
Support Computer (SSC) laptops (IBM 760XD ThinkPads #6071 & #6039)
with a new 3.0 Gb hard drive, and later the SAMS IUC (space
acceleration measurement system/interface control unit) laptop hard
drive.
Ed Lu also conducted another inventory audit (#4). Today's focus was
on battery packs for the CDMK and CSA-CP (carbon dioxide monitor kit
and compound specific analyzer-combustion products) instrumentation.
[Objective: To verify the stowage locations, current usage status,
and a count of the modified CDM/CSA-CP battery packs. With swapouts
not possible at present due to the Shuttle stand-down, the inventory
is required to determine the number of additional battery packs to be
manifested on 13P and subs. Referring to uplinked tabular lists, Ed
recorded location, status and configuration of the CHeCS (crew health
care systems) power supplies (there should be 31 packs onboard, 19
delivered on 9A and 12 on 11P).]
Malenchenko completed his regular daily inspection of the BIO-5
Rasteniya-2 ("Plants-2") experiment which studies growth and
development of plants (peas) under spaceflight conditions in the
Lada-2 greenhouse.
Both crewmembers conducted their regular daily physical exercise
program of 2.5 hrs., and Ed Lu completed the monthly maintenance of
the TVIS treadmill.
The Science Officer today performed the routine maintenance of SM
life support systems (SOZh) and the regular daily status checkup of
the autonomous Lab payloads, while the CDR prepared the daily
inventory management system (IMS) "delta" file for updating the IMS
database and completed..
Ed Lu began recording video scenes of a two-day on-board tour for a
"Live on ISS" PAO TV program, to be edited by him on 7/11 (Friday)
and later downlinked, with live crew narration, on 7/17 (Thursday
next week). At that time, the downlink will be aired on NASA TV.
[Scenes of interest include an end-to-end station "fly-through",
starting from the Lab and ending at the SM, highlights of the Lab
interior, the MSG (microgravity science glovebox), EarthKAM and TVIS
treadmill, the Airlock and its spacesuits, if visible, the FGB and
its stores, the Soyuz TMA-1 CRV, views of outside trusses, etc.]
EarthKAM operations are continuing on board, under remote commanding.
Today's CEO targets, no longer limited in the current LVLH attitude
and including the targets of the Lewis & Clark 200-year memorial
locations, were Eastern Mediterranean Aerosols (low visibility
reported between the Syrian coast and Cyprus [3-4 miles visibility],
probably smog from Italy. Looking left of track with sea as
background, and some coastal features to help orient interpreters.
Techniques for shooting aerosols are oblique views, shorter lenses,
and slight over-exposure if possible), Athens, Greece (nadir pass),
Madrid, Spain (nadir pass), Western Mediterranean Dust (low
visibility [1-4 miles] reported from cities in southernmost Spain and
along the Algerian coast. Smog expected as the culprit, probably
from the Rhone valley of France), Algiers, Algeria (nadir pass), Dar
es Salaam, Tanzania (nadir pass), Brasilia, Brazil (nadir pass), Las
Vegas, Nevada (looking a touch right), Tucson, AZ (looking right for
the city. Remnants of the big fire on the Catalina Mts just north of
town are still active), Machu Picchu, Peru (this famous site of Inca
ruins lies on a ridge in a great bend of the Urubamba River which can
be seen on both sides of this most visited tourist point in South
America), and Kinshasa, Zaire (nadir pass. Better opportunity than
usual). CEO images can be viewed at the websites
http://eol.jsc.nasa.gov
http://earthobservatory.nasa.gov
See also the website "Space Station Challenge" at
http://voyager.cet.edu/iss/
Besides the daily CEO (crew earth observation) target list, the
station residents were also provided with times of ISS equatorial
night crossings for observing the current equatorial-tropical storm
zone.
U.S. and Russian Segment Status (as of 1:30pm EST).
Environmental Control and Life Support (ECLSS) and Thermal Control
(TCS):
Elektron O2 generator is powered On (16 amp mode). Vozdukh CO2
scrubber is On (in Manual Mode 5/3). U.S. CDRA CO2 scrubber is Off.
TCCS (trace contaminant control subsystem) is operating. MCA (major
constituents analyzer) is operating. BMP Harmful Impurities unit:
absorbent bed #1 in Purify mode, bed #2 in Purify mode (unit is off).
RS air conditioner SKV-1 is On; SKV-2 is Off.
SM Working Compartment: Pressure (mmHg) -- 742; temperature (deg C)
-- 27.3; ppO2 (mmHg) -- data invalid; ppCO2 (mmHg) -- data invalid.
SM Transfer Compartment: Pressure (mmHg) -- 751; temperature (deg C)
-- 20.0.
FGB Cabin: Pressure (mmHg) -- 748; temperature (deg C) -- n/a.
Node: Pressure (mmHg) -- 740.07; temperature (deg C) -- 22.9 (shell);
ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
U.S. Lab: Pressure (mmHg) -- 743.39; temperature (deg C) -- n/a;
ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
Joint Airlock (Equip. Lock): Pressure (mmHg) -- 743.49; temperature
(deg C) -- n/a; shell heater temp (deg C) -- 23.3, ppO2 (mmHg) --
168.8; ppCO2 (mmHg) -- 3.9.
PMA-1: Shell heater temp (deg C) -- 23.3
PMA-2: Shell heater temp (deg C) -- 20.6.
(n/a = data not available)
Propulsion System (PS):
Total propellant load available: 3803 kg (8384 lb) as of 6/26
[SM(774) + FGB(2447) + Progress M(182) +Progress M-1(400)].
(Capability: SM -- 860 kg; FGB -- 6120 kg).
- Continued -
@Message posted automagically by IMTHINGS POST 1.30
---
* Origin: SpaceBase(tm) Pt 1 -14.4- Van BC Canada 604-473-9358 (1:153/719.1)SEEN-BY: 633/267 270 @PATH: 153/719 715 7715 140/1 106/2000 633/267 |
|
| SOURCE: echomail via fidonet.ozzmosis.com | |
Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.